# Weiqi Wang

### v1

@@ Abstract
To enable Large Language Models (LLMs) to function as conscious agents with generalizable reasoning capabilities, it is crucial that they possess the reasoning ability to comprehend situational changes (transitions) in distribution triggered by environmental factors or actions from other agents. Despite its fundamental significance, this ability remains underexplored due to the complexity of modeling infinite possible changes in an event and their associated distributions, coupled with the lack of benchmark data with situational transitions. Addressing these gaps, we propose a novel formulation of reasoning with distributional changes as a three-step discriminative process, termed as MetAphysical ReaSoning. We then introduce the first-ever benchmark, [Uncaptioned image]Mars, comprising three tasks corresponding to each step. These tasks systematically assess LLMs’ capabilities in reasoning the plausibility of (i) changes in actions, (ii) states caused by changed actions, and (iii) situational transitions driven by changes in action. Extensive evaluations with 20 (L)LMs of varying sizes and methods indicate that all three tasks in this process pose significant challenges, even for state-of-the-art LLMs and LMs after fine-tuning. Further analyses reveal potential causes for the underperformance of LLMs and demonstrate that pre-training them on large-scale conceptualization taxonomies can potentially enhance their metaphysical reasoning capabilities. Our data and models are publicly accessible at https://github.com/HKUST-KnowComp/MARS.
@@ 1 Introduction
Recent advancements in LLMs have demonstrated superior performance across a variety of reasoning tasks, such as causal reasoning [41, 12, 37], commonsense reasoning [51, 10, 31], mathematical reasoning [30, 21, 1], and more. However, to truly achieve conscious processing [3], the integration of System II reasoning ability [56, 35] is essential as it enables LLMs to perform out-of-distribution generalization when encountered with unfamiliar scenarios [8]. Among several components that make up System II reasoning, a critical element of it is the ability to reason with situational changes in distribution, triggered by environmental factors and actions by themselves or other agents, when dealing with non-stationarities [6]. It serves as the core ability in planning tasks [28], which can be achieved by dynamically recombining existing concepts in the given environment or action and learning from the resultant situational changes [38, 4, 17]. For instance, in the event that “PersonX is driving a car in a sunny day,” a change in the weather from sunny to rainy could cause a different outcome, such as “PersonX becomes more cautious and drives slower.” This illustrates that a change in weather conditions can lead to a change in the driver’s behavior, which represents an environmental change that triggers situational changes within the distribution of different weathers.
@@
Though fundamental, the exploration of this ability has been limited due to several factors. First, the scope for change within an event is vast, with numerous components capable of altering in a wide variety of ways. This results in an overwhelmingly large number of potential changes that are impossible to fully cover with existing knowledge bases. Second, reasoning with changes in distribution lacks a clear formulation due to its complexity. Unlike one-step inference reasoning tasks [54, 9], changes in action may lead to implausible events that cannot occur in reality, thus terminating the reasoning process. Lastly, there is a lack of a reliable evaluation benchmark. Existing benchmarks [62, 25] typically focus on a limited number of changes within a few scenarios, thus limiting the coverage of formed distributions. The changes in actions and states are also formulated under planning or logical tasks, which neglect transitions (consequences) caused by changes.
@@
To address these gaps, we take a step forward by formally defining reasoning with changes in distribution as a three-step discriminative process. We start by defining seven categories of changes, each corresponding to different components within an event. To semantically cover more changes in a unified manner, we propose implementing changes by altering each component within the event using their abstractions or numerical variations. This approach creates a hierarchical distribution of various changes, with the abstracted ones offering a more generalized coverage. Inspired by [8], we formulate reasoning with changes in distribution as sequentially tasking the model to: (1) assess the plausibility of a potential change in a given event that describes an action, (2) evaluate the plausibility of an inferential state resulting from the modified action, and (3) determine the necessary change in an action to convert an implausible inferential state into a plausible one. We refer to this process as metaphysical reasoning, as it requires models to distinguish implausible actions, states, and transitions that only exist in the metaphysical realm, indicating their rare occurrence in reality [26].
@@
We then construct the first evaluation benchmark, [Uncaptioned image]Mars, featuring 355K annotated data across three tasks corresponding to each step. It is constructed by sequentially instructing ChatGPT to extract events from Wikitext [46] and BookCorpus [79], identify mutable components within each event, generate abstractions and numerical variations for those components, create a metaphysical inference state based on the changes, and generate the necessary modifications to make the metaphysical inference plausible in reality. Large-scale human annotations are then conducted to verify the generations and construct them into an evaluation benchmark. Extensive experiments with over 20 language models demonstrate that all three tasks in this process present significant challenges, even for LMs after fine-tuning. Further analyses reveal potential reasons for such underperformance and identify possible solutions for enhancing the metaphysical reasoning abilities of language models.
In summary, the main contributions of our work are three-fold:
  • Task. We introduce the task of metaphysical reasoning, which includes three distinct subtasks, designed to assess the ability of (L)LMs to reason with changes in distribution.
  • Resource. We carefully curate a large-scale evaluation benchmark, [Uncaptioned image]Mars, to facilitate evaluations of LMs’ metaphysical reasoning abilities. The dataset is released.
  • Evaluation. We experiment with several (L)LMs to demonstrate the difficulty of our proposed tasks and conduct analysis to identify the reasons behind their underperformance.
@@ 2 Backgrounds and Related Works
Reasoning about Changes in Distribution. Enabling LMs to understand distributional changes due to localized causal interventions, particularly in semantic spaces, has long been a crucial objective in the pursuit of conscious machine intelligence [7, 8]. Previous works have mainly explored this within the context of discriminating changes between actions and states with methods such as commonsense knowledge injection [58], event calculus [5], and fuzzy reasoning [77]. Other studies aim to benchmark this reasoning process through logical reasoning tasks [25] and planning tasks [62, 71]. However, these studies only cover changes in limited formats and scenarios and also overlook the significance of representing them as a distribution in relation to different variables in actions. This restricts the out-of-distribution generalizability of the resulting LMs when facing unfamiliar scenarios. Moreover, previous evaluations do not cover transitions caused by changes, making subsequent evaluations incomplete. We address these issues by proposing to use the abstraction or numerical variations of components as changes to form generalizable distributions. We also design a task in Mars to evaluate LMs’ proficiency in understanding how changes motivate situational transitions.
Benchmarking LLMs. The advent of LLMs [48, 49, 61, 60, 53] has sparked various studies in investigating LLM’s potential in a variety of tasks, including text generation [45, 14, 13], temporal reasoning [57, 76], causal reasoning [12, 15, 34], commonsense reasoning [31, 10], and more [51]. These studies have significantly contributed to our understanding of LLMs by evaluating their performance across diverse tasks, using different scales of parameters and prompting methods [50]. However, there is an absence of a comprehensive benchmark for assessing the ability of (L)LMs to reason with changes in distribution. This inspires us to formally define it and introduce the first benchmark that evaluates such reasoning capabilities of (L)LMs.
@@ 3 Definitions of Changes in Event and Metaphysical Reasoning
Figure 1: Examples of changes in event in our formulation. Events may become metaphysical as components are abstracted into high-level concepts, while some remain plausible in reality.
Modeling changes within an event is inherently complex due to the infinite combinations that can occur. For simplicity, we only consider events that represent an action and changes between their inferential states. Given an event $e$, we first define seven types of changes that could transpire within $e$. These changes are represented as components of the event, including its subject $s$, verb $v$, object $o$, temporal quantifier $t$, spatial quantifier $l$, numerical properties $n$, and sub-events $se$. The original event is denoted as a function of these seven components, $e=f(s,v,o,t,l,n,se)$. A change in the event can be represented by altering one of its components, for instance, $e^{\prime}=f(s^{\prime},v,o,t,l,n,se)$ if the change impacts the subject $s^{\prime}$.
To effectively model the distribution of changes across different types of components, we leverage two types of hierarchical formulations. Specifically, for $s,v,o,se$, we define changes in these components as conceptualizing their original instance into three concepts with progressively increased abstractedness [22, 59]. For $t,l,n$, we define their changes as modifications from their original values to three distinct numerical or spatial values with progressively increased units. This brings a hierarchical structure to changes of a certain component, forming a distribution that gradually covers more possible changes. Abstracted conponents, as high-level concepts, can semantically represent a broader range of combinations for altering an event. Some running examples of how changes impact an action are shown in Figure 1.
We then propose a three-step discriminative process, which we term as Metaphysical Reasoning, to formulate reason with changes in distribution. The three steps, as shown in Figure 2, are:
@@
Figure 2: Three steps in metaphysical reasoning. Our motivation behind this is that, by conquering all steps sequentially, a conscious agent could answer: (1) Will the change occur in reality? (2) What will the change cause? (3) What change can make a metaphysical (desired) inference plausible?
@@
We then introduce our sequential pipeline for curating the [Uncaptioned image]Mars benchmark. An overview of our curation pipeline is shown in Figure 3. To guarantee a comprehensive coverage of events across various domains and topics in Mars, we source original text from two publicly available large corpora: Wikitext [46] and BookCorpus [79]. We filter out noisy text that includes hashtags, hyperlinks, and the like, and segment long text into sentences, each comprising no more than 200 tokens, to facilitate future processing.
4.1 Text Decomposition and Component Extraction
We first perform text decomposition [74, 32] to break down lengthy text into semantically complete short events, which are then used for fine-grained component extraction. To enable large-scale processing, we use ChatGPT [48], a powerful LLM with strong text understanding abilities, as the core processor for all stages. For each stage, we guide it with a few-shot prompt [69, 11] by creating task-specific explanations and exemplars (more details about prompts are in Appendix C):
@@ 4.1 Text Decomposition and Component Extraction
To perform text decomposition, <TASK-PROMPT> clarifies the goal to ChatGPT, which involves extracting semantically complete actions from the given text. <INPUT1-10> and <OUTPUT1-10> are filled with 10 pairs of human-crafted examples, each containing several action events extracted from text sampled from Wikitext and BookCorpus. ChatGPT is expected to learn from these examples and use them as a guide to extract action events (<OUTPUT(11,1-N)>) from the final input text (<INPUT11>).
For component extraction, we adjust <TASK-PROMPT> to define the task of extracting the seven components from a given event. We populate <INPUT1-10> and <OUTPUT1-10> with 10 pairs of events and seven comma-separated lists of components extracted from the event, each corresponding to one type of components defined in §3. ChatGPT then extracts seven lists of components for the final given event (<INPUT11>). If any type of component is absent, “None” will be generated instead.
@@ 4.2 Component Abstraction and Variation
The next step is designed to implement changes within the event by altering its components, extracted from the previous step, by generating their abstractions or numerical variations. Following [64], we guide ChatGPT by modifying <TASK-PROMPT> with the objective of generating abstract concepts for $s,v,o,se$ and numerical variations for $t,l,n$ within a specified event. For each <INPUT1-10> and <OUTPUT1-10> pair, we populate the input with a specific event and one of its components. The output consists of three human-authored component abstractions or numerical variations that align with the event’s context. Subsequently, ChatGPT is tasked with generating three abstractions or numerical variations for the final pair of the given event and a component within the event (<INPUT11>). Replacing the original components in the event with their generated changes forms changed event candidates for the metaphysical event discrimination task.
Figure 3: An overview of our benchmark curation pipeline with running examples.
@@ 4.5 Human Annotations
Annotation: Finally, we carry out large-scale human annotations to label candidate data for each task via Amazon Mechanical Turk (AMT). We provide detailed instructions with examples to qualified workers and task them with annotating (1) the plausibility of the changed events generated in §4.2, (2) the plausibility of the plausible/metaphysical inferences produced in §4.3, and (3) the plausibility of the transitions generated in §4.4. We collect five votes for each entry and the majority vote is used as the final label. The overall inter-annotator agreement (IAA) is 81% in terms of pairwise agreement, and the Fleiss kappa [20] is 0.56, indicating sufficient agreement. More details are in Appendix E.
Verification: To verify the quality of our collected labels, we recruit three postgraduate students with rich experience in NLP research to perform a second round annotation. Each of them is asked to annotate a sample of 100 data entries for each task, following the same instructions provided to the AMT annotators. The results are presented in Table 1. On average, 93.67% labels collected from human annotations align with the expert’s vote, demonstrating the reliability of our collected labels.
AbsATM [23] N/A 7,196 1.060 107,384 12,117 11,503 131,004 372,584 N/A
AbsPyramid [67] N/A 16,944 1.690 176,691 22,050 22,056 220,797 0 N/A
Meta. Event. 9,998 55,190 1.040 96,004 12,013 11,982 119,999 329,540 91.0%
AbsATM [23] N/A 7,196 6.413 65,386 8,403 7,408 81,197 5,921,195 N/A
Propara [16] 9,051 9,051 N/A 7,043 913 1,095 9,051 0 N/A
TRAC [25] 15,000 15,000 N/A 10,000 2,000 3,000 15,000 0 N/A
PlanBench [62] 26,250 26,250 N/A 0 0 26,250 26,250 0 N/A
Component Type Identified Modified
ME. MI. MT. #Avg. ME. MI. MT. #Avg.
Subject 4,376 3,907 3,507 1.116 3,106 2,950 2,591 1.094
Verb 9,874 8,856 8,061 3.647 4,408 4,146 3,760 3.457
Object 12,645 11,302 9,986 1.760 5,949 5,494 4,865 1.703
Temporal Quantifier 3,003 2,560 2,288 0.472 1,394 1,253 1,110 0.435
Spatial Quantifier 3,866 3,741 3,301 0.459 2,064 1,979 1,718 0.476
Numerical Properties 5,619 4,932 4,355 0.652 3,570 3,353 2,920 0.612
Sub-events 419 385 326 0.040 425 402 332 0.037
Total 39,802 35,683 31,824 8.146 20,916 19,577 17,296 7.814
5 Evaluations and Analysis
5.1 [Uncaptioned image]Mars Statistics
Table 1 presents statistics of the Mars benchmark, which comprises a total of 355,617 annotated data distributed across three tasks. We partition the annotated data into training, development, and testing splits following an 8:1:1 ratio, ensuring there is no overlap of text and events between the different splits to preserve the evaluation’s generalizability. On average, 1.4 tokens are
generated to describe changes in action for the metaphysical event and transition discrimination tasks, while 10.4 tokens are used for inferences in the metaphysical inference discrimination task. To the best of our knowledge, we are the first in proposing such a triad of tasks concurrently within a single benchmark. To compare Mars with other datasets, we select those with analogous task objectives for each task and compare them individually (see Appendix F). We find Mars tends to be significantly larger than other benchmarks, covering a broader range of events and providing training sets for evaluating the performance of fine-tuned models.
To further illustrate the diverse coverage of events and modified components in Mars, we match each component variation against hypernyms in Probase [72] and plot their distribution according to their number of occurrences in Figure 4. Our results indicate that Mars covers over 170,000 hypernyms in Probase, spanning broad categories such as event, activity, concept, unit, and more.
5.2 Main Evaluations on [Uncaptioned image]Mars
5.2.1 Task Setup and Model Selections
We then experiment with a selection of (L)LMs to investigate their performances on our curated Mars benchmark. Following the task definitions in §3, each task is assessed as a binary classification task. For unbiased evaluations, we use accuracy, ROC-AUC, and Macro-F1 scores as evaluation metrics.
@@ 5.2.1 Task Setup and Model Selections
Gemma-1.1-7B 56.88 - 48.53 51.83 - 51.76 49.41 - 45.01
@@
The evaluation of different models are categorized into three types: (1) Zero-shot: We first evaluate several (L)LMs in a zero-shot manner. For small-sized Pre-Trained Language Models (PTLMs), we evaluate RoBERTa [42], DeBERTa-v3 [24], GPT2 [52], CAR [63], CANDLE [64], and VERA [40], following the design of zero-shot question answering [44]. For LLMs, we evaluate LLaMa2, LLaMa3 [60, 61], Gemma [47], Falcon [2], and Mistral [33] using direct zero-shot prompting [51]. (2) Finetuning: We then assess the performance of (L)LMs when fine-tuned on the training set of Mars. For PTLMs, we fine-tune RoBERTa, DeBERTa, GPT2-xl, and VERA. For LLMs, we fine-tune LLaMa2, LLaMa3, Gemma, and Mistral using LoRA [27]. (3) LLM API: Finally, we evaluate the performance of ChatGPT [48] and GPT-4 [49], which represent proprietary LLMs, under zero-shot, five-shots, Chain-of-Thought prompting (COT; [68]), and Self-Consistent COT (SC-COT; [66]) settings. Please find implementation details in Appendix D, multi-task fine-tuning experiments in Appendix G, and few-shot fine-tuning experiments in Appendix H.
5.2.2 Results and Analysis
Evaluation results are reported in Table 3. From the results, we observe that:
(1) Most models exhibit subpar performance under the zero-shot setting. Among PTLMs, only VERA delivers acceptable results across all three tasks, while the rest significantly underperform. This indicates the extreme difficulty of our proposed metaphysical reasoning tasks. Though models fine-tuned on commonsense knowledge and conceptualizations [23, 65], such as CAR and CANDLE, show some improvement compared to their DeBERTa-v3-Large backbone, these performances are still unsatisfactory, even falling below the level of majority voting. For LLMs, LLaMa-3-70B outperforms all other LLMs on the three tasks, making it the best zero-shot model. Nevertheless, all models perform poorly across all tasks in Mars, emphasizing the difficulty of our tasks.
(2) Fine-tuning only offers limited benefits. With fine-tuning, all models improve significantly. For example, DeBERTa-Large’s accuracy increases by 16.18%, 21.84%, and 22.2% on three tasks, respectively. However, the best results for all tasks are still capped at around 74%, indicating a shared difficulty and significant room for future enhancements. One potential reason for this is that, since we split the data according to the source of text in Wikitext and BookCorpus, the distribution between different splits may differ significantly, as the domain and topics could be diverse from each other.
(3) The GPT series models underperform compared to other LLMs, and COT does not consistently aid performance. Surprisingly, GPT series models fall short when compared to open LLMs, such as LLaMa-3-70B. One possible explanation is that negative examples in Mars are sourced from ChatGPT’s generation and are obtained via post-human annotation. This makes it challenging to discriminate as these negative examples contradict ChatGPT’s internal knowledge. More advanced prompting methods, like COT, tend to negatively impact the models’ performance.

### v2

@@ Abstract
To enable Large Language Models (LLMs) to function as conscious agents with generalizable reasoning capabilities, it is crucial that they possess the ability to comprehend situational changes (transitions) in distribution triggered by environmental factors or actions from other agents. Despite its fundamental significance, this ability remains underexplored due to the complexity of modeling infinite possible changes in an event and their associated distributions, coupled with the lack of benchmark data with situational transitions. Addressing these gaps, we propose a novel formulation of reasoning with distributional changes as a three-step discriminative process, termed as MetAphysical ReaSoning. We then introduce the first-ever benchmark, [Uncaptioned image]Mars, comprising three tasks corresponding to each step. These tasks systematically assess LLMs’ capabilities in reasoning the plausibility of (i) changes in actions, (ii) states caused by changed actions, and (iii) situational transitions driven by changes in action. Extensive evaluations with 20 (L)LMs of varying sizes and methods indicate that all three tasks in this process pose significant challenges, even after fine-tuning. Further analyses reveal potential causes for the underperformance of LLMs and demonstrate that pre-training on large-scale conceptualization taxonomies can potentially enhance LMs’ metaphysical reasoning capabilities. Our data and models are publicly accessible at https://github.com/HKUST-KnowComp/MARS.
[Uncaptioned image]Mars: Benchmarking the Metaphysical Reasoning Abilities of Language Models with a Multi-task Evaluation Dataset
Weiqi Wang, Yangqiu Song Department of Computer Science and Engineering, HKUST, Hong Kong SAR, China {wwangbw, yqsong}@cse.ust.hk
@@ 1 Introduction
Figure 1: Examples of changes in event in our formulation. After changes occur, events may become metaphysical as components are abstracted into high-level concepts, while some remain plausible in reality.
@@
Recent advances in LLMs have demonstrated superior performance in a variety of reasoning tasks Liu et al. (2023b); Chan et al. (2024); Ko et al. (2023); Qin et al. (2023); Jain et al. (2023). However, to truly achieve conscious processing Andreas (2022), the integration of System II reasoning ability Sloman (1996); Kahneman (2011) is essential as it enables LLMs to perform out-of-distribution generalization when encountered with unfamiliar scenarios Bengio et al. (2021). Among several components that make up System II reasoning, a critical element of it is the ability to reason with situational changes in distribution, triggered by environmental factors and actions by themselves or other agents, when dealing with non-stationarities Bengio (2017). It serves as the core ability in planning tasks Huang et al. (2024), which can be achieved by dynamically recombining existing concepts in the given environment or action and learning from the resultant situational changes Lake and Baroni (2018); Bahdanau et al. (2019); de Vries et al. (2019). For instance, in the event that “PersonX is driving a car in a sunny day,” a change in the weather from sunny to rainy could cause a different outcome, such as “PersonX becomes more cautious and drives slower.” This illustrates that a change in weather conditions can lead to a change in the driver’s behavior, which represents an environmental change that triggers situational changes within the distribution of different weathers.
@@
Though fundamental, the exploration of this ability has been limited due to several factors. First, the scope for change within an event is vast, with numerous components capable of altering in a wide variety of ways. This results in an overwhelmingly large number of potential changes that are impossible to fully cover with existing knowledge bases. Second, reasoning with changes in distribution lacks a clear formulation due to its complexity. Unlike one-step inference reasoning tasks Sap et al. (2019), changes in action may lead to implausible events that cannot occur in reality, thus terminating the reasoning process. Such type of changes require extra care when designing evaluation protocols. Lastly, there is a lack of a reliable evaluation benchmark. Existing benchmarks Valmeekam et al. (2023); He et al. (2023b) typically focus on a limited number of changes within a few scenarios, thus limiting the coverage of formed distributions. The changes in actions and states are also formulated under planning or logical tasks, which neglect transitions (consequences) caused by changes.
@@
To address these gaps, we take a step forward by formally defining reasoning with changes in distribution as a three-step discriminative process. We start by defining seven categories of changes, each corresponding to different components within an event. To semantically cover more changes in a unified manner, we propose implementing changes by altering each component within the event using their abstractions or numerical variations. This approach creates a hierarchical distribution of various changes, with the abstracted ones offering a more generalized coverage. Inspired by Bengio et al. (2021), we formulate reasoning with changes in distribution as sequentially tasking the model to: (1) assess the plausibility of a potential change in a given event that describes an action, (2) evaluate the plausibility of an inferential state resulting from the modified action, and (3) determine the necessary change in an action to convert an implausible inferential state into a plausible one. We refer to this process as metaphysical reasoning–a term we adopt to describe a mode of reasoning that deals with highly improbable or abstract scenarios distinct from its traditional philosophical meaning or counterfactual reasoning (see Appendix A)–as it also requires models to distinguish implausible actions, states, and transitions that exist only in this abstract “metaphysical” realm, indicating their rare occurrence in reality Heidegger (2014).
@@
We then construct the first evaluation benchmark, [Uncaptioned image]Mars, featuring 355K annotated data across three tasks corresponding to each step. It is constructed by sequentially instructing an LLM to extract events from Wikitext Merity et al. (2017) and BookCorpus Zhu et al. (2015), identify mutable components within each event, generate abstractions and numerical variations for those components, create a metaphysical inference state based on the changes, and generate the necessary modifications to make the metaphysical inference plausible in reality. Large-scale human annotations are then conducted to provide labels of evaluation data entries and verify the quality of our benchmark. Extensive experiments with over 20 (L)LMs demonstrate that all three tasks in this process present significant challenges, even for LMs after fine-tuning. Further analyses reveal potential reasons for such underperformance and identify possible solutions for enhancing the metaphysical reasoning abilities of language models.
@@ 2 Backgrounds and Related Works
Reasoning about Changes in Distribution. Enabling LMs to understand distributional changes due to localized causal interventions, particularly in semantic spaces, has long been a crucial objective in the pursuit of conscious machine intelligence Bengio et al. (2019, 2021). Previous works have mainly explored this within the context of discriminating changes between actions and states with methods such as commonsense knowledge injection Tandon et al. (2018), event calculus Basina et al. (2022), and fuzzy reasoning Zhang et al. (2013). Other studies aim to benchmark this reasoning process through logical reasoning tasks He et al. (2023b) and planning tasks Valmeekam et al. (2023); Wu et al. (2021). However, these studies only cover changes in limited formats and scenarios and also overlook the significance of representing changes as a distribution in relation to different variables in actions. Such loss restricts the out-of-distribution generalizability of the resulting LMs when facing unfamiliar scenarios. Moreover, previous evaluations do not cover transitions caused by changes, making subsequent evaluations around reasoning with changes incomplete.
Figure 2: The three steps in metaphysical reasoning. Our motivation behind this is that, by conquering all steps sequentially, a conscious agent could answer: (1) Will the change occur in reality? (2) What will the change cause? (3) What change can make a metaphysical (desired) inference plausible?
Benchmarking LLMs. The advent of LLMs OpenAI (2022, 2023); Touvron et al. (2023b, a); Reid et al. (2024) has sparked various studies in investigating LLM’s potential in a variety of tasks Chen et al. (2024b, a); Yuan et al. (2024); Chan et al. (2024); Jain et al. (2023); Qin et al. (2023). These studies have significantly contributed to our understanding of LLMs by evaluating their performance across diverse tasks, using different scales of parameters and prompting methods Qiao et al. (2023). However, there is an absence of a comprehensive benchmark for assessing the ability of (L)LMs to reason with changes in distribution. This inspires us to formally define it and introduce the first benchmark that evaluates such reasoning capabilities of (L)LMs.
@@ 3 Definitions of Changes in Event and Metaphysical Reasoning
Modeling changes within an event is inherently complex due to the infinite number of changes that can occur. For simplicity, we only consider events that represent an action and study changes between their inferential states. Given an event $e$, we first define seven types of changes that could transpire within $e$. These changes are represented as components of the event, including its subject $s$, verb $v$, object $o$, temporal quantifier $t$, spatial quantifier $l$, numerical properties $n$, and sub-events $se$. The original event is denoted as a function of these seven components, $e=f(s,v,o,t,l,n,se)$. A change in the event can be represented by altering one of its components, for instance, $e^{\prime}=f(s^{\prime},v,o,t,l,n,se)$ if the change impacts the subject $s^{\prime}$.
To effectively model the distribution of changes across different types of components, we leverage two types of hierarchical formulations. Specifically, for $s,v,o,se$, we define changes in these components as conceptualizing their original instance into three concepts with progressively increased abstractedness Giunchiglia and Walsh (1992); Tenenbaum et al. (2011). For $t,l,n$, we define their changes as modifications from their original values to three distinct numerical or spatial values with progressively increased units. This brings a hierarchical structure to changes of a certain component, forming a distribution that gradually covers more possible changes. Abstracted components, as high-level concepts, can semantically represent a broader range of combinations for altering an event. Some running examples of how changes impact an action are shown in Figure 1. We then propose a three-step discriminative process, which we term as Metaphysical Reasoning (see Appendix A), to formulate reason with changes in distribution. The three steps, as shown in Figure 2, are:
@@
We then introduce our sequential pipeline for curating the [Uncaptioned image]Mars benchmark. An overview of our curation pipeline is shown in Appendix Figure 5. To guarantee a comprehensive coverage of events across various domains and topics, we source original text from two publicly available large corpora: Wikitext Merity et al. (2017) and BookCorpus Zhu et al. (2015). We filter out noisy text that includes hashtags and hyperlinks and segment long text into sentences with no more than 200 tokens to facilitate future processing.
4.1 Text Decomposition and Extraction
We first perform text decomposition Ye et al. (2023); Jhamtani et al. (2023) to break down lengthy text into semantically complete short events, which are then used for fine-grained component extraction. To enable large-scale processing, we use ChatGPT OpenAI (2022), a powerful LLM with strong text understanding abilities, as the core processor for all stages. For each stage, we guide it with a few-shot prompt West et al. (2022); Brown et al. (2020) by creating task-specific explanations and exemplars (detailed prompts are in Appendix B):
@@ 4.1 Text Decomposition and Extraction
To perform text decomposition, <TASK-PROMPT> clarifies the goal to ChatGPT, which involves extracting semantically complete actions from the given text. <INPUT1-10> and <OUTPUT1-10> are filled with 10 pairs of human-crafted examples, each containing several action events extracted from text sampled from Wikitext and BookCorpus. ChatGPT is expected to learn from these examples and use them as a guide to extract action events (<OUTPUT(11,1-N)>) from the final input text (<INPUT11>). For component extraction, we adjust <TASK-PROMPT> to define the task of extracting the seven components from a given event. We populate <INPUT1-10> and <OUTPUT1-10> with 10 pairs of events and seven comma-separated lists of components extracted from the event, each corresponding to one type of components defined in §3. ChatGPT then extracts seven lists of components for the final given event (<INPUT11>). If any type of component is absent, “None” will be generated instead.
@@ 4.2 Component Abstraction and Variation
The next step is designed to implement changes within the event by altering its components, extracted from the previous step, by generating their abstractions or numerical variations. Following Wang et al. (2024b), we guide ChatGPT by modifying <TASK-PROMPT> with the objective of generating abstract concepts for $s,v,o,se$ and numerical variations for $t,l,n$ within a specified event. For each <INPUT1-10> and <OUTPUT1-10> pair, we populate the input with a specific event and one of its components. The output consists of three human-authored component abstractions or numerical variations that align with the event’s context. Subsequently, ChatGPT is tasked with generating three abstractions or numerical variations for the final pair of the given event and a component within the event (<INPUT11>). Replacing the original components in the event with their generated changes forms changed event candidates for the metaphysical event discrimination task.
AbsATM He et al. (2024) N/A 7,196 1.060 107,384 12,117 11,503 131,004 372,584 N/A
AbsPyramid Wang et al. (2024d) N/A 16,944 1.690 176,691 22,050 22,056 220,797 0 N/A
Meta. Event. 9,998 55,190 1.040 96,004 12,013 11,982 119,999 329,540 94.0%
AbsATM He et al. (2024) N/A 7,196 6.413 65,386 8,403 7,408 81,197 5,921,195 N/A
Propara Dalvi et al. (2018) 9,051 9,051 N/A 7,043 913 1,095 9,051 0 N/A
TRAC He et al. (2023b) 15,000 15,000 N/A 10,000 2,000 3,000 15,000 0 N/A
PlanBench Valmeekam et al. (2023) 26,250 26,250 N/A 0 0 26,250 26,250 0 N/A
@@ 4.5 Human Annotations
Annotation: Finally, we carry out large-scale human annotations to label candidate data for each task via Amazon Mechanical Turk (AMT). We provide detailed instructions with examples to qualified workers and task them with annotating (1) the plausibility of the changed events generated in §4.2, (2) the plausibility of the plausible/metaphysical inferences produced in §4.3, and (3) the plausibility of the transitions generated in §4.4. We collect five votes for each entry and the majority vote is used as the final label. The overall inter-annotator agreement (IAA) is 81% in terms of pairwise agreement, and the Fleiss Kappa Fleiss (1971) is 0.56, indicating sufficient agreement (see Appendix D).
Expert Verification: To verify the quality of our collected labels, we recruit three postgraduate students with rich experience in NLP to perform a second round annotation. Each of them is asked to annotate a sample of 100 data entries for each task, following the same instructions provided to the AMT annotators. Results in Table 1 show that, on average, 93.67% labels collected from human annotations align with the expert’s vote, demonstrating the reliability of our collected labels.
@@
Meta-LLaMa-3.1-8B 51.01 - 50.27 52.13 - 51.29 52.35 - 52.09
Meta-LLaMa-3.1-70B 59.22 - 52.08 63.61 - 61.90 61.28 - 61.03
   +RAG 61.21 - 54.51 66.38 - 65.90 61.53 - 61.22
   +Multi-Agent 56.12 - 51.08 65.06 - 65.01 62.54 - 62.19
   +Self-reflection 57.94 - 53.17 63.91 - 63.51 60.92 - 60.77
Meta-LLaMa-3.1-405B 60.01 - 52.99 64.52 - 63.23 61.74 - 61.76
Gemma-2-9B 56.88 - 48.53 51.83 - 51.76 49.41 - 45.01
@@
GPT-4o-mini 57.94 - 57.91 53.84 - 53.53 48.06 - 48.06
   +RAG 59.99 - 59.97 54.54 - 54.21 49.39 - 49.19
   +Multi-Agent 54.21 - 53.17 52.76 - 52.26 46.94 - 46.70
   +Self-reflection 56.89 - 55.21 53.22 - 53.20 48.51 - 48.45
5 Evaluations and Analysis
5.1 [Uncaptioned image]Mars Statistics
Table 1 presents statistics of the Mars benchmark, which comprises a total of 355,617 annotated data distributed across three tasks. We partition the annotated data into training, development, and testing splits following an 8:1:1 ratio, ensuring there is no overlap of text and events between the different splits to preserve the evaluation’s generalizability. On average, 1.04 tokens are generated to describe changes in action for the metaphysical event and transition discrimination tasks, while 10.4 tokens are used for inferences in the metaphysical inference discrimination task. To the best of our knowledge, we are the first in proposing such a triad of tasks concurrently within a single benchmark. To compare Mars with other datasets, we select those with analogous task objectives for each task and compare them individually. We find Mars tends to be significantly larger than other benchmarks, covering a broader range of events and providing training sets for evaluating the performance of fine-tuned models. To further illustrate the diverse coverage of events and changes in Mars, we match each component variation against hypernyms in Probase Wu et al. (2012) and plot their distribution according to their number of occurrences in Figure 3. Our results indicate that Mars covers over 170,000 hypernyms in Probase, spanning broad categories such as event, activity, concept, unit, etc.